what is the equation of a line that passes through the points (5, 6) and (1, 3)
Answer by mananth(16949)

x1 y1 x2 y2
5 6 1 3

slope m = (y2-y1)/(x2-x1)
( 3 - 6 )/( 1 - 5 )
( -3 / -4 )
m= 3/4
The required line passes through both the points. So we select one point and plug it with the slope
Plug value of the slope and point( 1 , 3 ) in
Y = m x + b
3.00 = 3/4 * 1 + b

b= 3 - 3/4
b= 9/ 4
So the equation will be
Y = 3/4 x + 9/4
m.ananth@hotmail.ca